OpenStudy (anonymous):

0.02x-0.03y=0.16 0.07x+0.08y=0.92 solve for x and y

OpenStudy (anonymous):

Are you allowed to use technology to get your answer? This one is kinda a pain by hand.

OpenStudy (anonymous):

I just used technology to get the answer and it is REALLY ugly. You do NOT want to do that one by hand if you can avoid it.

OpenStudy (texaschic101):

.02x - .03y = 0.16 .07x + .08y = 0.92 multiply them both by 100 to get rid of the decimals giving us : 2x + 3y = 16 7x + 8y = 92 does that help ?

OpenStudy (anonymous):

No response means they don't really want help..just answers.

OpenStudy (ybarrap):

$$ 0.02x-0.03y=0.16\\ 0.07x+0.08y=0.92 $$ Multiply everything by 100: $$ 2x-3y=16\\ 7x+8y=92 $$ Multiply the 1st equation by 7 and the second by -2 $$ 7(2x-3y)=(16)(7)\\ -2(7x+8y)=(92)(-2) $$ Add the equations: $$ -21y-16y=16\times 7-184=112-184=-72\\ 37y=72\\ \color{red}{\mathbf{ y=\cfrac{72}{37}}\\} 2x-3\cfrac{72}{37}=16\\\ \color{red}{\mathbf{x={16+3\cfrac{72}{37}\over 2}}} $$